OpenStudy (anonymous):

A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days: f(n) = 10(1.02)n Part A: When the scientist concluded his study, the height of the plant was approximately 11.04 cm. What is a reasonable domain to plot the growth function? (4 points) Part B: What does the y-intercept of the graph of the function f(n) represent? (2 points) Part C: What is the average rate of change of the function f(n) from n = 1 to n = 5, and what does it represent? (4 points)

OpenStudy (anonymous):

@misssunshinexxoxo

OpenStudy (misssunshinexxoxo):

Are you familiar with x and y intercepts?

OpenStudy (anonymous):

yes cutie

OpenStudy (misssunshinexxoxo):

What is the height of the plant? "the height of the plant was approximately 11.04 cm"

OpenStudy (misssunshinexxoxo):

That's your y intercept

OpenStudy (anonymous):

yes so do we plug that in the equation that they give us

OpenStudy (anonymous):

and ok

OpenStudy (misssunshinexxoxo):

f(1) = 10.02 f(5) = 11.041 m(ave) = (11.041 - 10.02) / (5 - 1) = 0.255 cm/day

OpenStudy (anonymous):

so that whole thing would be part A

OpenStudy (misssunshinexxoxo):

It's all three

OpenStudy (misssunshinexxoxo):

Domain is 0 through 5

OpenStudy (misssunshinexxoxo):

What is the height of the plant?

OpenStudy (anonymous):

oh ok thank you your the best <3

OpenStudy (anonymous):

11.04cm

OpenStudy (misssunshinexxoxo):

That's B^

OpenStudy (misssunshinexxoxo):

y intercept

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

hey for part B what would the f(n) represent

OpenStudy (anonymous):

@misssunshinexxoxo

OpenStudy (misssunshinexxoxo):

The starting height of the plant

OpenStudy (anonymous):

oh ok that makes sense thank you

OpenStudy (anonymous):

so for the final part the function is going to look like this f(n)=10(1.02)^1 f(n)=10(1.02)^5

OpenStudy (anonymous):

@misssunshinexxoxo